Gregorian Technical Collimation Updates - 19 October 2011 UT

Observer(s): D. Thompson, K. Boutsia, A. Rakich, also present: A.!DiPaola, F.Pedichini, M.Centrone
Telescope Operator: G. (LBTO)
Telescope Support: K.Newton (LBTO)
Instrument Support: D. Officer (LBTO)

Summary

Details

01:30 Open dome

01:54 Problems with setting the shell - call JC Guerra

02:00 pointing correction

02:02 Active preset to BS9166
  • problems with the collimation
  • it did not recognize the X astigmatism and did not correct it at all - could it be because of the hot spot?
  • to fix this we need to go to higher elevation

02:16 Active preset to WT10_321
  • preset failed - home the AGW

02:50 AGWs restarted - new pointing model in place - new GCS configuration also (restart GCS)
  • pointing correction
  • shell ripped while we were pointing WT10_321 in acquire mode

03:02 Active preset to WT10_321 - GS=1
  • Hotspot is (x,y): 274.00, 298.00 - current position
  • new hotspot positions -> Hotspot is (x,y): 266.00, 288.00
  • restart GCS

03:28 Active preset to WT10_321 - GS=1 - new hotspot positions + binnng 2!!
  • the wfsc

03:36 field_collect, BS9107_field_0.list
  • elevation 25
  • it failed reaching the second star - pointing correction
  • agw reset - restarting GCS

04:04 field_collect, BS9107_field_0.list, START=2

04:11 we changed the second GS to 38 because the 45 seems to be too far out.
  • field_collect, BS9107_field_0, START=2
  • preset failed again - agw reset (the y was negative)

04:19 field_collect, BS9107_field_0, START=2
  • we lost 2 images of the off-axis collimation
  • we need to repeat

04:28 field_collect, BS9107_field_0, START=2
  • agw failure again!!!!
  • changing to MODS

04:31 start reconfiguring for MODS
  • problems with modsWake
  • MODS is not waking up - connected to the power loss this afternoon(?)

05:37 Going up the dome to check if MODS electronic boxes have power.

06:22 Reconfiguring to LUCI because we are not able to turn on MODS

06:52 pointing correction

07:00 field_collect, BS9107_field_0
  • start over, do both stars
  • it does the first star correctly but at the second GS
  • it does not find the *fits files in Repository/UT/GCS - maybe it takes too long to transfer images...

07:19 changing star again for the second iteration to GS=35
  • field_collect, BS9107_field_0, START=2
  • we exited before finishing collimation on-axis
  • andrew has a log of the images
GStar wfsc #
0 144-146
35 147-149
0 150-152

07:27 field_collect, BS9107_field_0
  • we need collimation data on the first star again...
GStar wfsc #
0 153-154
36 155-157
0 158-159
  • implementing modifications in binodal astigmatism

07:37 field_collect, BS9107_field_0
  • repeat field_collect to see if the modification is working
GStar wfsc #
0 164-165
36 166-168
0 169
35 170-172
0 173-175

07:55 we need to do a ptmodify - after andrew's calculations

07:58 field_collect, BS9107_field_0
GStar wfsc #
0 187-188
36 189-191
0 192-193
  • failed going to the second star - agw restart

08:12 field_collect, BS9107_field_0, START=2
  • preset failed - same AGw problem
  • leave the PC off for 10 min.

08:33 restarting AGw computer
  • pointing correction

08:40 field_collect, BS9107_field_0, START=2
GStar wfsc #
0 197-198
35 199-201
0 202
  • the changes in M2 X Y RX RY for correcting binodal astigmatism are the following global offsets:
    dx=0.255, dy=0.766, rx=-147.3, ry=49.1
These changes correspond to a tip of -15 arc seconds and a tilt of -5 arc seconds. If the set of field data we are taking now confirms these measurements, then we will also want to modify IE and CA by adding 15 arc seconds to IE and adding 5 arc seconds to CA in the new pointing model. To be confirmed first....

08:48 field_collect, BS9107_field
  • failed during the 4th off-axis star because of the same AGW problem
  • we switch to the pisces focal station

09:04

10:50 Start collecting DX RFBG collimation data

Had to modify rangebal to make M2max=7.0 for the DX M2.

hsel @wfs.lis "M1COLL0,M1COLL1,M1COLL2,M1COLL3,M1COLL4,M2COLL0,M2COLL1,M2COLL2,M2COLL3,M2COLL4,TELEL,T301,T201,AMBTEMP,GLASTEMP,$I" yes > "hdr.dat"

right_wfscimage000035.fits 87 WT10_316
right_wfscimage000126.fits 83 WT10_301
right_wfscimage000042.fits 82 WT10_299
right_wfscimage000047.fits 77 ACT0299
right_wfscimage000132.fits 76 ACT0316
right_wfscimage000052.fits 72 WT10_267
right_wfscimage000140.fits 67 WT10_344
right_wfscimage000021.fits 64 ACT0231
right_wfscimage000119.fits 64 WT10_233
right_wfscimage000057.fits 61 WT10_231
right_wfscimage000145.fits 55 WT10_329
right_wfscimage000114.fits 54 WT10_197
right_wfscimage000065.fits 49 ACT0178
right_wfscimage000152.fits 46 WT10_342
right_wfscimage000107.fits 44 WT10_161
right_wfscimage000071.fits 40 WT10_142
right_wfscimage000096.fits 35 WT10_127
right_wfscimage000079.fits 30 WT10_110
right_wfscimage000087.fits 26 BS9123

File edited to include {}s and descriptive info.

-- KonstantinaBoutsia - 19 Oct 2011
Topic revision: r6 - 19 Oct 2011, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback